OpenStudy (anonymous):

Can anybody point me to online resources that talk about using de moivre's theorem to prove trig identities like below: sin 2x = 2 cos x. sin x

OpenStudy (anonymous):

hmm. the idea is this. \[e^{i\theta}=cos(\theta)+isin(\theta)\] and \[(e^i\theta)^2=e^{2i\theta}=(cos(\theta)+isin(\theta))^2=cos({2\theta})+isin(2\theta)\]

OpenStudy (anonymous):

square \[(cos(\theta) + i sin(\theta))\] to see what you get. then equate the real part to the real part and you get an identity for \[cos(2\theta)\] and another one for \[sin(2\theta)\]

OpenStudy (anonymous):

when you compute \[(cos(\theta)+isin(\theta))^2\] you get \[cos^2(\theta)-sin^2(\theta)+i\times 2 cos(\theta)sin(\theta)\] the real part is \[cos^2(\theta)-sin^2(\theta)\] so that must equal the real part of \[cos(2\theta)+isin(2\theta)\] which is just \[cos(2\theta)\] telling you that \[cosd(2\theta)=cos^2(\theta)-sin^2(\theta)\]

OpenStudy (anonymous):

likewise \[sin(2\theta)=2cos(\theta)sin(\theta)\]

OpenStudy (anonymous):

My concern is: what de moivre's theorem says? z^n = r^n (( cos nx) + i sin (n x)) or (cos x + i sin x) ^n= ( cos nx) + i sin (n x) I am confused.

OpenStudy (anonymous):

both are true.

OpenStudy (anonymous):

How come> can u explain?

OpenStudy (anonymous):

\[z=re^{i\theta}=r(cos(\theta)+isin(\theta))\]

OpenStudy (anonymous):

\[z^n=r^n(e^{i\theta})^n=r^ne^{ni\theta}\]

OpenStudy (anonymous):

that by the laws of exponents.

OpenStudy (anonymous):

and since \[e^{ni\theta}=cos(n\theta)+isin(n\theta) \] you get the second equality

OpenStudy (anonymous):

if you have not seen \[z=re^{i\theta}\] as a representation of a complex number, then it requires a difffernt explanation, but if you have seen it it is nothing more than the laws of exponents.

OpenStudy (anonymous):

succinctly put here http://en.wikipedia.org/wiki/De_Moivre%27s_formula

OpenStudy (anonymous):

Thanks .I need to site on this, it has been itching my head since yesterday.I appreciate your help.

OpenStudy (anonymous):

welcome hope at least second explanation was clear.

OpenStudy (anonymous):

i must have made another algebra error let me check.

OpenStudy (anonymous):

if you have not seen \[z=re^{i\theta}\] as a representation of a complex number, then it requires a difffernt explanation, but if you have seen it it is nothing more than the laws of exponents.